OpenStudy (anonymous):

add /simplify if possible 7/9z^3 + 1/21z^2

jhonyy9 (jhonyy9):

7/9z3 + 1/21z2 =z2(7/9z +1/21)=z2(7*7/63z +1*3/63)=z2(49z +3)/63